OpenStudy (hannaha):

MEDAL FOR HELP!! PLEASE HELP EXPLAIN THIS!! AS IN HOW TO DO IT--NOT JUST THE ANSWER!! Simplify this expression. Type the variable term before the constant. 12 + 3(2x - 3) + 4

OpenStudy (ineedhelplz):

Let's simplify step-by-step. 12+3(2x−3)+4 Distribute: =12+(3)(2x)+(3)(−3)+4 =12+6x+−9+4 Combine Like Terms: =12+6x+−9+4 =(6x)+(12+−9+4) =6x+7

OpenStudy (mrhirohito):

This topic is really part of studying the Order of Operations, but simplifying with parentheses is probably the sub-topic that causes students the most difficulty, so this lesson is meant to provide a little extra help in this area. When simplifying expressions with parentheses, you will be applying the Distributive Property. That is, you will be distributing over (multiplying through) a set of parentheses in order to simplify a given expression. I will walk you through some examples of increasing difficulty, and you should note, as this lesson progresses, the importance of simplifying as you go and of doing each step neatly, completely, and exactly.

OpenStudy (mrhirohito):

That is just a little help! ineed is the best person to have with you ^_^

OpenStudy (hannaha):

thanks both of you :)

OpenStudy (hannaha):

@Ineedhelplz So basically you're multiplying 3x2 and 3x3?

OpenStudy (hannaha):

Then you just combine the like terms (6x by itself) and do the math for the remaining numbers? (12 + -9 + 4= 7) is this right?